OpenStudy (anonymous):

Given f ^{2}=g _{2}+h _{2} where g(x) and h(x) are polynomials with real coefficients, prove that f has non-real roots

myininaya (myininaya):

\[ f ^{2}=g _{2}+h _{2} \]?

myininaya (myininaya):

are g and h suppose to be squared too?

myininaya (myininaya):

or did you mean for them to be subscripts?

myininaya (myininaya):

i think you meant \[f^2=g^2+h^2\]

myininaya (myininaya):

because you gave us no info for the functions g_2 and h_2

OpenStudy (anonymous):

yeah, g and h are supposed to be square...

OpenStudy (anonymous):

The solution that you want to show is not true. Consider \[g(x)=h(x)=x\] Then \[f(x)^2=2x^2\] so \[f(x)=\pm\sqrt{2}x\] which has no non-real roots.

OpenStudy (anonymous):

However, if you state that at least one of g or h has a nonzero constant term and that they are not both constant functions, then clearly \[f^2(x) > 0\] for all real x, and since f^2 is clearly a polynomial, it must have at least one non-real root by the fundamental theorem of algebra